OpenStudy (anonymous):

Solve the equation (eqn coming soon) . . . . ANSWER = {1/4,4}

OpenStudy (anonymous):

\[(\frac{ x+2 }{ x-1 })^2+(\frac{ x+2 }{ x-1 })-6=0\]

hartnn (hartnn):

let y= x+2 / x-1

OpenStudy (anonymous):

any hints?? how should i start

OpenStudy (anonymous):

ok

hartnn (hartnn):

u get quadratic in y

OpenStudy (anonymous):

ok i am solving

OpenStudy (anonymous):

woah! it was so easy the look of the question was confusing

OpenStudy (anonymous):

not really!

hartnn (hartnn):

y= ?

OpenStudy (anonymous):

\[y=\frac{ -1\pm \sqrt{-23} }{ 2 }\]

OpenStudy (anonymous):

now what?

hartnn (hartnn):

i m getting 2,-3

OpenStudy (anonymous):

Me too.

OpenStudy (anonymous):

;( how?

OpenStudy (anonymous):

You might have done a mistake..

OpenStudy (anonymous):

y^2+y-6=0 is the equation right??

hartnn (hartnn):

yes

hartnn (hartnn):

b^2-4ac = ?

hartnn (hartnn):

1-(-24) = ?

OpenStudy (anonymous):

oh c was -6 i did it +6

OpenStudy (anonymous):

now i have to put the values of y right??

hartnn (hartnn):

yes x+2 / x-1 = 2 x+2 / x-1 =-3 cross multiply

OpenStudy (anonymous):

first one x=4

hartnn (hartnn):

:)

OpenStudy (anonymous):

I got only one and it's x = 4

OpenStudy (anonymous):

the second is x=1/4

hartnn (hartnn):

\(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

@Miyuru we have to put y 2 times as we got it 2 "2 and 3"

OpenStudy (anonymous):

oops 2 and -3

OpenStudy (anonymous):

Ohh yeah I forgot... :D

OpenStudy (anonymous):

Did you checked your answer in the first equation... :)